OpenStudy (anonymous):

what is the domain of p? p(x)2- 2x+3

OpenStudy (anonymous):

x is 2 to the square just don't know how to write it

OpenStudy (anonymous):

p=x^2-2x+3?

OpenStudy (anonymous):

p(x)=x^2-2x+3. is that it??

OpenStudy (anonymous):

In that case the domain is the entire real axis. You can plug any value in for x as there are no roots or divisions.

OpenStudy (anonymous):

thank u for showing me where that sign was Thomas. u are a life saver

OpenStudy (anonymous):

If the above equation is written correctly, then @Thomas9 is right. You start off with \[(-\infty,\infty)\]and then from there you look for any square roots or division by a variable to start eliminating points from the domain, but in this case you have neither. So we are just left with what we started with, \[(-\infty,\infty)\]

OpenStudy (anonymous):

thank u